OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. Choose the ordered pair that is a solution to this system of equations: 3x + y = 5 4x + 5y = 3 (1, -1) (1, -2) (1, 2) (2, -1) (2, 1)

OpenStudy (anonymous):

these questions are all over the map is this some sort of on line course?

OpenStudy (anonymous):

answer is x = 2, y = -1 but i am not sure how this relates to previous questions

OpenStudy (experimentx):

3x + y = 5 4x + 5y = 3 y = 5 - 3x 4x + 5(5 - 3x )= 3 x = 2, 2*2 + y = 5 y = -1 c